40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1953     Entry time: Mon Jun 29 16:59:47 2015
Author: Arjun 
Type: Misc 
Category: PD noise 
Subject: Pre-stabilisation of laser  

We have managed to pre stabilise the laser using a few SR560's. It is not as stable as the one that would be implemented once the digital system is in place. But it should be good enough for some preliminary data. As described in a previous log we had locked the cavity to track the laser. The PZT actuator in the cavity is driven by a SR560 which has a limited output voltage range from (-4V to +4V) and if due to slow frequency drifts the frequency of the laser drifts beyond the limit to which the PZT can compensate the cavity would unlock itself. Also, currently the intensity noise of laser has not been stabilised, this passes directly through the cavity and will appear at the transmission if its not accounted for. This feedback has been achieved using a AOM. The AOM is driven by a RF function generator and the output power of thr RF function generator can be modulated by this in turn changes the power in the carrier frequency by pushing some power into the first order diffracted beam, thereby stabilising the laser intensity fluctuation by almost 2 orders of magnitude. A brief description of the feed-back loops is given below.

Frequency Stabilisation

The laser was frequency stabilised for its slow drifts, this could lead to SR650 controlling the cavity not being able to compesnsate for it due to its limited output range. The output of cavity stabilisation was low passed and then fed to the frequency control of the laser. But the gain had to be adjusted appropriately, as otherwise the loop could become unstable. This was done by using a simple resistive divider circuit (potentiometer) to first attenuate the control signal for the cavity stabilisation and then low passed and fed back to counter for the slow frequency drifts of the laser.The cutoff freqeucncy of the secondary feedback loop is also important, so as to ensure that at cross-over frequency nyquist stability condition is satisfied. The cutoff for this was kept at 30mHz.

Intensity stabilisation 

Intensity fluctuations from the laser will appear at the output port of the PMC because of the fact that the cavity is locking itself to laser's frequency. These have to be corrected for in the final setup and this has been achieved using an AOM. A AOM splits power in the beam into a diffracted beam and this splitting of power depends on the power that is injected into the AOM and also the direction in whihc the AOM is oriented. When no power is injected in the AOM the carrier beam passes through unaffected and when some non-zero power is injected a part of power goes into another diffracted beam. The AOM we use has a maximum power input of 2W at 80MHz. For  achieving the required functionality, a RF signal from a RF signal generator is amplified using a high power RF amplifier and this drives the AOM. Now to stabilise the intensity fluctuations going into the PMC we can setup a positive feedback by sensing the power at thePMC's output and using that to modulate the signal generated by RF signal generator thereby modulating the power with which the AOM is driven and finally controlling the way power is split between the two output rays. Hence this way the power entering the PMC has been stabilised. In the actual setup this feedback will be provided by the common mode output of the two photodiodes being tested.The entire pre-stabilisation of laser implemented is shown in the schematic below.

The total stabilisation schematic is shown below-this includes the locking of the cavity, feedback to supress the laser frequency drift and supression of intensity fluctuations.

We were also able to get some more optics fixed, since the table will also be used for another experiment, we divided the beam using a halfwave plate and a polarising beam splitter to control the power going into each of the experiments. A image of the setup assembled is attached below, the read trace is the path of the laser.

Attachment 1: Screenshot_2015-06-29-16-20-55.png  421 kB  | Hide | Hide all
Screenshot_2015-06-29-16-20-55.png
Attachment 6: 20150626_202413.jpg  2.168 MB  | Hide | Hide all
20150626_202413.jpg
ELOG V3.1.3-